OpenStudy (onthecly2):

Find the quotient of the complex numbers: 2 * (sqrt 3) (cos (116 deg) + i sin (116 deg)) / 2(cos (82 deg) + isin (82 deg ))

OpenStudy (irishboy123):

\[\dfrac{ 2 * (sqrt 3) (cos (116 deg) + i sin (116 deg)) }{ 2(cos (82 deg) + isin (82 deg ))} \] that?! [i won't tighten the Latex]

OpenStudy (onthecly2):

The problem literally has a division symbol instead of a fraction set up so I don't know...

OpenStudy (holsteremission):

If \(z=re^{i\theta}\) and \(w=se^{i\varphi}\), then \[z/w=\frac{z}{w}=\frac{r}{s}e^{i(\theta-\varphi)}\]i.e. divide the moduli and subtract the arguments.

OpenStudy (irishboy123):

if there is a problem with a division vs fraction symbol, i reckon the idea of polar complex is not on the menu, hence my laziness in the latex hope i am wrong :)

OpenStudy (sshayer):

\[\frac{ 2\sqrt{3} e^{\iota 116}}{ 2 e^{\iota82} }=\sqrt{3}e ^{^{\iota(116-82)}}=\sqrt{3}e ^{\iota34}\] =?

OpenStudy (sshayer):

change in cos and sin

OpenStudy (irishboy123):

previous one on OS:: http://openstudy.com/users/onthecly2#/updates/58519d6de4b02201566815f1